OpenStudy (anonymous):

need help asap!!!!! For each sale, you should fill in the information in the appropriate cell in the chart. When you are finished, submit this worksheet to your teacher following the directions indicated in the lesson. Information Sales Discount Given as a Percentage Off Discount Given as a Percentage Off Discount Given as Item Marked Down Discount Given as Item Marked Down Place, date of discovery, and wording of the sign Proposed item to be purchased Origina

OpenStudy (anonymous):

theres no info on what to work out

OpenStudy (anonymous):

There is no data. :/

OpenStudy (anonymous):

Oh.. I get it, you have to find some things on sale.

OpenStudy (anonymous):

Have you found any items yet?

OpenStudy (texaschic101):

Information : Jewelry marked $143 discount given as percent off --- 30% final price (without tax) = 143 - 143(.30) = 143 - 42.90 = $100.10 place -- Wal-Mart date of discovery -- pick a date wording on sign -- 30% off all jewelry on this table Information : All clothing marked down 50% discount given as percent off --- 50% prices before discount vary place -- JC Penny date of discovery -- pick a date wording on sign -- 50% off all clothing on this rack information : sale on fans discount given on items marked down -- 35.99 marked down to 25 place -- Best Buy date of discovery -- pick a date wording on sign -- great sale on fans information : CD's marked down discount given on items marked down -- 12.99 marked down to 10 place -- Hastings date of discovery -- pick a date wording on sign -- CD's marked down

OpenStudy (texaschic101):

I am sorry but I have to go now